# MILU: A Multi-task Indic Language Understanding Benchmark

### v2

@@ Abstract
Evaluating Large Language Models (LLMs) in low-resource and linguistically diverse languages remains a significant challenge in NLP, particularly for languages using non-Latin scripts like those spoken in India. Existing benchmarks predominantly focus on English, leaving substantial gaps in assessing LLM capabilities in these languages. We introduce MILUMulti-task Indic Language Understanding Benchmark—a comprehensive evaluation benchmark designed to address this gap. MILU spans 8 domains and 42 subjects across 11 Indic languages, reflecting both general and culturally specific knowledge. With an India-centric design, MILU incorporates material from regional and state-level examinations, covering topics such as local history, arts, festivals, and laws, alongside standard subjects like science and mathematics. We evaluate over 45 LLMs, and find that current LLMs struggle with MILU, with GPT-4o achieving the highest average accuracy at 72%. Open multilingual models outperform language-specific fine-tuned models, which perform only slightly better than random baselines. Models also perform better in high-resource languages as compared to low-resource ones. Domain-wise analysis indicates that models perform poorly in culturally relevant areas like Arts & Humanities and Law & Governance compared to general fields like STEM. To the best of our knowledge, MILU is the first of its kind benchmark focused on Indic languages, serving as a crucial step towards comprehensive cultural evaluation. All code, benchmarks, and artifacts will be made publicly available to foster open research.
@@ 1 Introduction
In this work, we introduce MILU-Multi-task Indic Language Understanding Benchmark, a comprehensive evaluation dataset designed to address these gaps. MILU spans 8 domains and 42 subjects across 11 Indic languages, reflecting both general and culturally specific knowledge. We designed MILU with an India-first perspective by collecting questions from various national, state, and regional exams. These questions include culturally relevant subjects such as local history, arts, festivals, and laws, alongside traditional academic subjects like science and mathematics. Following previous efforts Hendrycks et al. (2021); Zhong et al. (2023), we create this benchmark by collecting questions from over 1500 competitive exams from India. We focus on region-specific exams to authentically capture local knowledge in the respective language.
We evaluate 45 different LLMs - a mix of closed proprietary, open-source, and language-specific models- on MILU. Our findings suggest that models struggle with MILU, with GPT-4o achieving the highest average accuracy at 72%. Interestingly, open multilingual models outperform language-specific models, which only achieve slightly better than random scores. Our analysis of in-context learning reveals that adding more examples improves performance in base models, but the effect on instruct models remains inconclusive. We also explore how performance scales with the number of parameters, finding significant improvements as model size increases. Our domain-wise analysis reveals that models perform poorly in culturally relevant areas, such as Arts & Humanities and Social Sciences, compared to more general fields like STEM. All the artifacts will be released publicly.
@@ 3.1 Questions Curation
MILU is a large, multi-domain test set containing multiple-choice based questions (MCQs) taken from over 42 subjects with an emphasis on India-specific knowledge. This benchmark covers many domains, including Mathematics, Science, Social Sciences, Humanities, Arts, Business Studies, and Law, among others. MILU is designed as a culturally relevant benchmark to assess general problem-solving abilities and India-specific knowledge. These questions were sourced following an approach similar to AGIEval Zhong et al. (2023), collecting the questions from various public exams taken by individuals intending to either pursue higher studies or seek career advancements, such as qualification tests and national and state-level civil services exams, among others.
@@
In total, we collected more than 150K questions across 11 Indian Languages- Bengali (bn), Gujarati (gu), Hindi (hi), Kannada (kn), Malayalam (ml), Marathi (mr), Odia (or), Punjabi (pn), Tamil (ta), Telugu (te), and English (en)-spanning 42 diverse subjects. English questions are also included as these often address Indian culture-specific content, which is notably missing from existing popular benchmarks. Table 1 presents a brief comparison of the type of questions present in MMLU Hendrycks et al. (2021) and MILU.
@@ 3.2 Data Cleaning and Filtering
Upon examination, we found that approximately 45% of questions were accurately labeled with a topic name, while the remaining questions lacked this information. To address this issue, we first translate the untagged questions into English using IndicTrans2 Gala et al. (2023) and then prompt GPT-4o-mini model to assign an appropriate topic name to the question. Finally, in total, we get around 20K tags. However, these tags are highly fine-grained, often having a heavy overlap. To organize them, we embed the tags using the NV-Embed-v2 Lee et al. (2024) model and apply K-means clustering to group tags into 50 clusters. We manually review these clusters and assign appropriate subject labels. Following the manual merging of related clusters, we determine 42 distinct subject names, which fall into eight main domains: Arts and Humanities, Social Sciences, Environmental Sciences, Law and Governance, Health and Medicine, Mathematics and Sciences, Engineering and Technology, and Business Studies. The final distribution of domains within MILU is shown in Figure 2. Detailed analysis of question distribution by topic & language is provided in Appendix B.
@@
In total, we release around 85K questions across 42 subjects across 8 domains in 11 languages, capping each subject-language pair at 500 questions for feasible evaluations.
@@ 3.3 Data Analysis
Table 2 shows the overall statistics of MILU. Of the total 85K questions, only 18% of questions are translated from English, with the remainder preserved in their original source languages. The average question length varies across languages, with languages such as kn, ml, and ta having shorter word counts due to their agglutinative nature.
@@
Science & Math: Includes topics in physics, chemistry, biology, and mathematics, with references to Vedic Math, Ayurveda, and ancient as well as modern scientific achievements.
@@ 5 Results and Discussions
GPT-4o 75.32 72.06 77.16 74.85 70.96 71.79 67.68 71.85 67.16 70.51 80.66 72.73
GPT-4o-mini 59.13 55.94 62.04 55.59 53.27 54.66 46.12 56.7 49.49 52.04 69.4 55.85
Gemini-1.5-Pro 68.11 66.93 69.93 68.83 64.95 65.58 61.68 64.58 65.13 63.75 76.46 66.90
Gemini-1.5-Flash 60.71 60.31 63.86 63.25 58.65 58.71 55.98 61.12 57.06 56.82 68.53 60.45
Krutrim-spectre-v2 42.74 41.61 44.5 44.17 40.66 39.44 35.18 39.81 38.82 38.17 45.83 40.99
@@
meta-llama/Llama-3.2-1B 25.28 25.49 26.90 25.24 25.05 26.46 24.93 24.82 25.72 25.75 34.08 26.34
meta-llama/Llama-3.2-1B-Instruct 25.79 25.71 25.31 25.85 25.09 25.56 24.53 25.02 26.51 26.09 25.79 25.57
sarvamai/sarvam-2b-v0.5 29.83 29.06 31.15 28.69 25.87 28.77 26.85 28.72 27.78 28.30 34.47 29.04
sarvamai/sarvam-1 32.99 33.18 32.90 32.98 30.34 31.43 30.48 32.22 30.01 31.69 35.63 32.17
google/gemma-2-2b 29.77 28.89 32.26 29.67 29.05 30.98 25.63 30.48 30.06 28.43 50.18 31.40
google/gemma-2-2b-it 30.58 29.34 34.16 28.70 29.16 32.03 26.80 28.99 30.40 29.29 51.85 31.94
meta-llama/Llama-3.2-3B 32.74 30.94 34.27 31.41 28.97 31.94 27.98 30.30 30.54 29.69 49.59 32.58
meta-llama/Llama-3.2-3B-Instruct 32.97 32.77 30.42 31.61 30.10 28.35 26.64 32.34 30.99 31.17 31.68 30.82
nvidia/Nemotron-4-Mini-Hindi-4B-Base 31.33 25.83 50.53 27.53 27.10 38.60 26.68 26.38 27.96 28.17 50.54 32.79
@@
Telugu-LLM-Labs/Navarasa-2.0 36.01 36.88 39.51 39.42 36.05 36.61 35.00 36.35 36.13 36.25 51.04 38.11
neulab/Pangea-7B 42.14 37.11 41.97 34.18 34.62 36.98 32.89 36.35 32.93 36.17 55.87 38.29
CohereForAI/aya-23-8B 28.04 27.01 35.96 27.39 28.45 30.67 27.20 26.99 28.91 26.30 45.14 30.19
meta-llama/Llama-3.1-8B 37.60 35.44 42.29 36.87 34.53 37.51 32.63 36.16 35.57 34.63 58.83 38.37
meta-llama/Llama-3.1-8B-Instruct 34.19 33.30 35.37 35.69 34.00 32.26 29.91 32.89 33.65 33.08 45.66 34.55
google/gemma-2-9b 52.60 50.61 55.52 53.28 49.50 49.39 40.01 48.81 48.81 47.99 68.87 51.40
google/gemma-2-9b-it 46.07 39.02 44.49 43.31 41.11 38.02 34.68 40.29 38.68 38.91 63.08 42.51
google/gemma-2-27b 62.86 58.96 63.28 62.54 59.35 59.17 45.11 59.15 57.27 57.00 74.00 59.88
google/gemma-2-27b-it 59.47 55.04 60.49 57.55 53.36 54.43 45.65 54.84 54.37 53.57 72.90 56.52
@@
CohereForAI/aya-23-35B 32.83 30.05 42.96 31.37 32.35 36.71 31.14 27.50 34.62 29.19 52.27 34.64
meta-llama/Llama-3.1-70B 50.53 47.09 47.51 49.01 47.94 46.88 47.7 47.55 47.66 46.33 55.33 48.5
meta-llama/Llama-3.1-70B-Instruct 61.59 58.66 57.92 61.50 57.83 53.05 52.00 59.54 56.45 56.26 53.41 57.11
meta-llama/Llama-3.1-405B 64.31 59.77 50.36 62.14 62.20 49.70 53.97 60.06 57.20 58.40 60.86 58.09
Table 3: Evaluation results of all the Closed and Open Multilingual models supporting all languages on MILU. We report 5-shot accuracies for all open models (except for Llama-3.1-70B, Llama-3.1-70B-Instruct & Llama-3.1-405B for which we report 0-shot accuracy) with the accuracy averaged across all the domains per language. Higher values indicate better model performance for the given language. Refer Section (§5.1) for more details.
@@ 5.1 How do models supporting all languages perform?
Table 3 shows the performance of various models supporting all the languages in the 5-shot setup. Among these, GPT-4o emerges as the strongest model, consistently outperforming all competitors across languages, with an average accuracy of 72.73%. Closed proprietary models, generally achieve higher performance than the open-source alternatives. Among the open-source models, Llama-3.1-70B shows the best performance.
@@ 5.2 How do language-specific fine-tuned models perform?
hi OpenHathi-7B-Hi 29.74 29.46 28.69 27.75 31.52 28.60 27.56 30.58 29.24
Airavata 27.97 29.03 29.89 27.15 27.44 28.10 26.91 33.00 28.69
ProjectIndus 26.19 25.69 26.89 25.04 23.05 25.40 25.44 25.56 25.41
Gajendra-v0.1 29.30 29.15 25.65 29.30 27.21 29.34 28.38 29.47 28.48
open-aditi-v6-llama3 29.29 32.74 31.02 30.15 33.96 32.42 31.96 33.58 31.89
AryaBhatta-GemmaGenZ-Merged 39.39 39.15 38.48 41.14 39.28 40.26 35.87 48.08 40.21
te telugu-llama-7b-instruct-v0.1 32.12 26.67 26.93 24.23 25.34 25.93 25.63 24.52 26.42
Telugu-Llama2-7B-v0-Base 22.17 27.43 25.25 27.77 25.53 26.87 25.32 27.45 25.97
Telugu-Llama2-7B-v0-Instruct 25.30 21.07 24.99 23.54 23.62 26.43 27.55 27.86 25.05
ka Kan-LLaMA-7B-base 28.88 29.19 30.50 27.35 28.30 26.49 29.25 30.50 28.81
Ambari-7B-base-v0.1 31.49 27.08 29.26 28.62 27.82 27.88 27.94 31.00 28.89
Ambari-7B-Instruct-v0.1 28.26 27.91 24.92 26.90 23.54 28.46 26.95 27.00 26.74
mr Misal-7B-base-v0.1 27.09 26.56 25.28 26.70 26.36 25.59 26.79 23.89 26.03
Misal-7B-instruct-v0.1 24.82 20.95 21.51 22.69 20.62 23.78 23.06 22.10 22.44
ta tamil-llama-7b-instruct-v0.2 22.17 23.53 25.03 24.64 28.53 26.29 25.29 22.57 24.76
ml malayalam-llama-7b-instruct-v0.1 30.00 28.44 23.43 27.87 25.00 26.37 23.33 20.00 25.56
@@ 5.5 How do the models perform on different domains?
hi OpenHathi Y N -0.36% -0.28% 0.59%
Airavata Y Y 1.54% 2.14% 2.93%
Gajendra-v0.1 Y Y 0.65% 0.85% 3.14%
kn Ambari-base Y N 0.75% 0.86% 2.07%
Ambari-Instruct Y Y 0.70% 0.46% 1.84%
Kan-7B-base Y N 0.76% 1.85% 2.11%
te telugu-llama-instruct Y Y 0.83% 0.56% 1.31%
Telugu-7B-v0-Base Y N 0.43% -0.79% 0.81%
Telugu-7B-v0-Instruct Y Y -1.02% -0.16% 0.32%
mr Misal-7B-base-v0.1 Y N 0.07% -0.44% -1.11%
Misal-7B-instruct-v0.1 Y Y -2.45% -2.04% -2.43%
ta tamil-llama-instruct Y Y -0.06% 0.22% 0.08%
ml malayalam-llama-instruct Y Y 0.62% 1.05% 0.45%

### v3

@@ Abstract
Evaluating Large Language Models (LLMs) in low-resource and linguistically diverse languages remains a significant challenge in NLP, particularly for languages using non-Latin scripts like those spoken in India. Existing benchmarks predominantly focus on English, leaving substantial gaps in assessing LLM capabilities in these languages. We introduce MILUMulti-task Indic Language Understanding Benchmark—a comprehensive evaluation benchmark designed to address this gap. MILU spans 8 domains and 41 subjects across 11 Indic languages, reflecting both general and culturally specific knowledge. With an India-centric design, MILU incorporates material from regional and state-level examinations, covering topics such as local history, arts, festivals, and laws, alongside standard subjects like science. We evaluate over 42 LLMs, and find that current LLMs struggle with MILU, with GPT-4o achieving the highest average accuracy at 74%. Open multilingual models outperform language-specific fine-tuned models, which perform only slightly better than random baselines. Models also perform better in high-resource languages as compared to low-resource ones. Domain-wise analysis indicates that models perform poorly in culturally relevant areas like Arts & Humanities and Law & Governance compared to general fields like STEM. To the best of our knowledge, MILU is the first of its kind benchmark focused on Indic languages, serving as a crucial step towards comprehensive cultural evaluation. All code, benchmarks, and artifacts will be made publicly available to foster open research.
@@ 1 Introduction
In this work, we introduce MILU-Multi-task Indic Language Understanding Benchmark, a comprehensive evaluation dataset designed to address these gaps. MILU spans 8 domains and 41 subjects across 11 Indic languages, reflecting both general and culturally specific knowledge. We designed MILU with an India-first perspective by collecting questions from various national, state, and regional exams. These questions include culturally relevant subjects such as local history, arts, festivals, and laws, alongside traditional academic subjects like science. Following previous efforts Hendrycks et al. (2021); Zhong et al. (2023), we create this benchmark by collecting questions from over 1500 competitive exams from India. We focus on region-specific exams to authentically capture local knowledge in the respective language.
We evaluate 45 different LLMs - a mix of closed proprietary, open-source, and language-specific models- on MILU. Our findings suggest that models struggle with MILU, with GPT-4o achieving the highest average accuracy at 74%. Interestingly, open multilingual models outperform language-specific models, which only achieve slightly better than random scores. Our analysis of in-context learning reveals that adding more examples improves performance in base models, but the effect on instruct models remains inconclusive. We also explore how performance scales with the number of parameters, finding significant improvements as model size increases. Our domain-wise analysis reveals that models perform poorly in culturally relevant areas, such as Arts & Humanities and Social Sciences, compared to more general fields like STEM. All the artifacts will be released publicly.
@@ 3.1 Questions Curation
MILU is a large, multi-domain test set containing multiple-choice based questions (MCQs) taken from over 41 subjects with an emphasis on India-specific knowledge. This benchmark covers many domains, including Science, Social Sciences, Humanities, Arts, Business Studies, and Law, among others. MILU is designed as a culturally relevant benchmark to assess general problem-solving abilities and India-specific knowledge. These questions were sourced following an approach similar to AGIEval Zhong et al. (2023), collecting the questions from various public exams taken by individuals intending to either pursue higher studies or seek career advancements, such as qualification tests and national and state-level civil services exams, among others.
@@
In total, we collected more than 150K questions across 11 Indian Languages- Bengali (bn), Gujarati (gu), Hindi (hi), Kannada (kn), Malayalam (ml), Marathi (mr), Odia (or), Punjabi (pn), Tamil (ta), Telugu (te), and English (en)-spanning 41 diverse subjects. English questions are also included as these often address Indian culture-specific content, which is notably missing from existing popular benchmarks. Table 1 presents a brief comparison of the type of questions present in MMLU Hendrycks et al. (2021) and MILU.
@@ 3.2 Data Cleaning and Filtering
Upon examination, we found that approximately 45% of questions were accurately labeled with a topic name, while the remaining questions lacked this information. To address this issue, we first translate the untagged questions into English using IndicTrans2 Gala et al. (2023) and then prompt GPT-4o-mini model to assign an appropriate topic name to the question. Finally, in total, we get around 20K tags. However, these tags are highly fine-grained, often having a heavy overlap. To organize them, we embed the tags using the NV-Embed-v2 Lee et al. (2024) model and apply K-means clustering to group tags into 50 clusters. We manually review these clusters and assign appropriate subject labels. Following the manual merging of related clusters, we determine 41 distinct subject names, which fall into eight main domains: Arts and Humanities, Social Sciences, Environmental Sciences, Law and Governance, Health and Medicine, Science, Engineering and Technology, and Business Studies. The final distribution of domains within MILU is shown in Figure 2. Detailed analysis of question distribution by topic & language is provided in Appendix B.
@@
In total, we release around 79K questions across 41 subjects across 8 domains in 11 languages, capping each subject-language pair at 500 questions for feasible evaluations.
@@ 3.3 Data Analysis
Table 2 shows the overall statistics of MILU. Of the total 79K questions, only 25% of questions are translated from English, with the remainder preserved in their original source languages. The average question length varies across languages, with languages such as kn, ml, and ta having shorter word counts due to their agglutinative nature.
@@
Science: Includes topics in physics, chemistry, biology, with references to Ayurveda, and ancient as well as modern scientific achievements.
@@ 5 Results and Discussions
GPT-4o 77.59 81.75 74.64 78.62 76.93 72.57 73.44 70.07 73.84 70.15 72.53 74.74
GPT-4o-mini 60.69 70.52 57.84 63.14 56.92 54.52 55.76 47.31 57.89 50.84 53.31 57.16
Gemini-1.5-Pro 68.8 76.91 69.68 70.63 69.68 65.5 65.63 62.21 65.37 66.03 64.47 67.72
Gemini-1.5-Flash 61.93 69.42 61.52 64.81 64.22 59.56 59.44 56.88 62.23 58.89 58.13 61.55
Krutrim-spectre-v2 44.48 46.59 43.43 45.39 46.05 41.89 40.89 36.8 40.57 40.48 39.22 42.34
@@
meta-llama/Llama-3.2-1B 25.37 34.35 25.46 27.01 25.51 25.27 26.5 25.02 24.52 25.56 25.79 26.4
meta-llama/Llama-3.2-1B-Instruct 25.17 25.84 25.19 25.24 25.91 25.02 25.82 24.4 24.74 26.8 26.11 25.48
sarvamai/sarvam-2b-v0.5 29.83 34.73 29.5 31.37 29.03 25.94 28.96 27.09 29.06 27.68 28.57 29.25
sarvamai/sarvam-1 33.31 30.3 29.83 33.05 33.29 30.8 27.9 29.33 28.76 27.64 28.5 30.25
google/gemma-2-2b 29.72 50.86 28.82 32.32 29.77 28.72 30.99 25.28 30.64 30.1 28.46 31.43
google/gemma-2-2b-it 30.6 52.4 29.25 34.33 28.52 28.81 32.18 26.32 29.01 30.46 29.22 31.92
meta-llama/Llama-3.2-3B 32.75 50.25 31.32 34.63 31.75 29.07 32.41 28.07 30.42 30.93 29.75 32.85
meta-llama/Llama-3.2-3B-Instruct 32.9 31.74 32.82 30.55 31.83 30.36 28.54 26.48 32.35 31.32 31.19 30.92
nvidia/Nemotron-4-Mini-Hindi-4B-Base 31.41 51.37 25.85 51.42 27.4 27.01 39.37 26.9 26.47 27.89 28.25 33.03
@@
Telugu-LLM-Labs/Navarasa-2.0 36.43 51.69 37.44 39.86 40.01 36.40 37.00 35.36 36.86 37.18 36.86 38.64
neulab/Pangea-7B 42.06 56.41 36.81 42.21 33.83 33.86 36.73 32.18 36.45 32.72 36.43 38.15
CohereForAI/aya-23-8B 27.85 45.76 26.93 36.31 27.53 28.4 30.96 27.51 26.98 28.88 26.34 30.31
meta-llama/Llama-3.1-8B 37.89 59.66 35.26 42.69 37.09 34.71 37.65 32.86 36.33 36.25 34.94 38.67
meta-llama/Llama-3.1-8B-Instruct 34.53 46.22 33.77 35.73 36.04 34.37 33.28 30.08 33.33 34.15 33.6 35.01
google/gemma-2-9b 53.12 69.8 51.19 56.24 53.99 49.55 49.81 40.42 49.7 50.05 48.78 52.06
google/gemma-2-9b-it 46.04 63.61 38.82 44.82 43.47 40.8 37.62 34.19 40.42 38.83 39.44 42.55
google/gemma-2-27b 63.47 74.82 59.29 64.33 63.43 59.62 59.66 44.86 59.94 58.6 57.94 60.54
google/gemma-2-27b-it 59.82 73.7 55.17 61.25 58.23 53.37 54.78 44.99 55.55 55.6 54.29 56.98
@@
CohereForAI/aya-23-35B 32.9 53.09 30.35 43.53 31.39 32.42 37.22 31.27 27.27 35 29.34 34.89
meta-llama/Llama-3.1-70B 67.37 74.59 62.42 66.97 68.37 62.69 62.2 57.48 64.6 60.7 60.64 64.37
meta-llama/Llama-3.1-70B-Instruct 62.77 54.00 60.26 58.90 63.02 58.94 53.71 52.81 60.67 58.41 57.66 58.29
meta-llama/Llama-3.1-405B 65.15 61.42 61.11 50.98 63.69 63.64 50.21 55.35 61.84 58.81 59.67 59.26
Table 3: Evaluation results of all the Closed and Open Multilingual models supporting all languages on MILU. We report 5-shot accuracies for all open models (except for Llama-3.1-70B-Instruct & Llama-3.1-405B for which we report 0-shot accuracy) with the accuracy averaged across all the domains per language. Higher values indicate better model performance for the given language. Refer Section (§5.1) for more details.
@@ 5.1 How do models supporting all languages perform?
Table 3 shows the performance of various models supporting all the languages in the 5-shot setup. Among these, GPT-4o emerges as the strongest model, consistently outperforming all competitors across languages, with an average accuracy of 74.74%. Closed proprietary models, generally achieve higher performance than the open-source alternatives. Among the open-source models, Llama-3.1-70B shows the best performance.
@@ 5.2 How do language-specific fine-tuned models perform?
hi OpenHathi-7B-Hi 29.72 28.59 28.81 27.74 30.96 28.53 28.28 30.76 29.17
Airavata 28.03 28.45 28.92 27.14 27.59 28.95 27.54 32.77 28.67
ProjectIndus 26.21 25.83 26.95 25.03 22.93 24.91 25.13 24.41 25.18
Gajendra-v0.1 29.30 28.36 25.54 29.30 27.75 29.18 28.38 30.26 28.51
open-aditi-v6-llama3 29.30 31.65 29.99 30.15 34.78 32.57 32.43 37.12 32.25
AryaBhatta-GemmaGenZ-Merged 39.42 37.88 37.59 41.14 39.60 41.25 35.71 48.49 40.14
te telugu-llama-7b-instruct-v0.1 30.19 27.20 26.27 24.03 27.16 26.41 26.81 24.63 26.59
Telugu-Llama2-7B-v0-Base 22.96 27.10 24.80 27.01 26.14 27.13 25.42 27.53 26.01
Telugu-Llama2-7B-v0-Instruct 26.03 24.00 24.45 23.68 24.48 26.82 27.17 28.01 25.58
ka Kan-LLaMA-7B-base 29.00 27.42 30.37 27.42 28.47 26.92 29.28 30.50 28.67
Ambari-7B-base-v0.1 31.49 27.52 29.08 28.70 27.15 27.82 27.93 31.00 28.84
Ambari-7B-Instruct-v0.1 28.17 28.02 24.83 27.78 23.84 27.56 27.42 27.00 26.83
mr Misal-7B-base-v0.1 26.96 26.65 25.02 27.19 27.28 26.11 26.82 24.05 26.26
Misal-7B-instruct-v0.1 25.05 21.44 21.55 22.89 21.82 24.59 23.36 22.16 22.86
ta tamil-llama-7b-instruct-v0.2 22.22 24.12 25.32 24.89 28.69 24.98 25.07 22.79 24.76
ml malayalam-llama-7b-instruct-v0.1 30.00 27.92 23.42 27.88 25.00 27.57 23.33 20.00 25.64
@@ 5.5 How do the models perform on different domains?
hi OpenHathi Y N -0.28% -0.25% 0.58%
Airavata Y Y 1.70% 2.40% 3.11%
Gajendra-v0.1 Y Y 0.73% 0.92% 3.22%
kn Ambari-base Y N 0.72% 1.17% 2.46%
Ambari-Instruct Y Y 0.60% 0.49% 1.87%
Kan-7B-base Y N 1.13% 0.84% 1.30%
te telugu-llama-instruct Y Y 1.13% 0.84% 1.30%
Telugu-7B-v0-Base Y N 0.32% -0.52% 0.88%
Telugu-7B-v0-Instruct Y Y -0.98% -0.12% 0.29%
mr Misal-7B-base-v0.1 Y N 0.16% -0.60% -0.93%
Misal-7B-instruct-v0.1 Y Y -2.55% -2.17% -2.38%
ta tamil-llama-instruct Y Y -0.06% -0.04% -0.19%
ml malayalam-llama-instruct Y Y 0.77% 1.09% 0.33%